Mean =12 hours ,a sample has 100 dry cells test to find the length of life , standard deviation ,= 3 hours assuming the data to be normally distributed, what percentage of battery is expected to have life more than 14 hrs , less than 18 hrs , between 10 and 14 hrs
1
Expert's answer
2021-07-26T14:24:45-0400

"N=100 \\

\mu=12 \\

\sigma=3 \\

P(X>14) = 1 -P(X< 14) \\

= 1 -P(Z< \frac{14-12}{3}) \\

= 1 -P(Z< 0.6666) \\

= 1 - 0.7476 \\

= 0.2524 \\

= 25.24 \; \% \\

P(X<18) = P(Z< \frac{18-12}{3}) \\

= P(Z< 2) \\

= 0.9772 \\

= 97.72 \; \% \\

P(10<X<14) = P(X<14) -P(X<10) \\

= P(Z< \frac{14-12}{3}) -P(Z< \frac{10-12}{3}) \\

= P(Z< 0.6666) -P(Z< -0.6666) \\

= 0.7476 -0.2527 \\

= 0.49 \\

= 49 \%"